OpenStudy (anonymous):

integrate: (0 to 5) for y(sqrt(25 − (y^2)) − 5 + y)dy should i take u substitution? for u=25 − (y^2), but how do i get rid of the other y?

OpenStudy (anonymous):

there is a y outside y[(sqrt(25 − (y^2)) − 5 + y)]dy

OpenStudy (lalaly):

\[\int\limits_{0}^{5}(y \sqrt{25-y^2-5+y})dy\]is this it?

OpenStudy (zarkon):

I think it might be \[\int\limits_{0}^{5}y\left[\sqrt{25-y^2}-5+y\right]dy\]

OpenStudy (anonymous):

zarkon got it

OpenStudy (zarkon):

distribute your outside y...then do two integrals...one with substitution and one without

OpenStudy (anonymous):

hmm will try ...